1. Multi-millionaire lottery Suppose you bought a lottery ticket at $100 today, January 5, 2018. You will find out the outcome on January 5, 2019. Let X be the possible dollar amount of the payoff you may receive on January 5, 2019 from the lottery. Assume the prevailing interest rate over the next 10 years is flat at 5% (per year). The likelihood of the possible outcome is assumed as follows. 4 Prob(X) 11/369/36 7/36 5/36 3/36 1/36 (1) What is your expected payoff on January 5, 2019? (2) What are the variance /standard deviation of your expectation in (1)? (3) What is vour exnected gain or loss on January 5. 2019?

(4) Suppose that because of federal regulations, you cant receive the payoff immediately after you won the lottery but 5 years later (on January 5, 2024), you will receive it with interest payments. Then, what is your compounded expected payoff on January 5, 2024? (5) Is there any correlation between the payoff amount X and the probability associated with it? (6) If we run the regression analysis using the amount X as its independent variable and the probability as its dependent variable, what would be the sign of regression coefficient?

(7) What is the meaning of the regression coefficient? (8) How can we tell whether the regression coefficients are statistically significant or not? Hint: Name the test.
Answer #1

Que 1

Expected payoff is the sum of probability adjusted payoff across all the possible outcomes = Xe =

sum_{i=1}^{6} P_{i} imes X_{i} where Pi is the probability of payoff Xi

= 1 x 11/36 + 2 x 9/36 + 3 x 7/36 + 4 x 5/36 + 5 x 3/36 + 6 x 1 / 36 = $ 91 / 36 mn = $ 2.53 mn

Que 2

Variance of discrete distribution, sigma2

= sum_{i=1}^{6} P_{i} imes (X_{i} - X_{e})^{2} where Xe = expected payoff as calculated in Que 1 above

= (1 - 2.53)2 x 11/36 + (2 - 2.53)2 x 9/36 + (3 - 2.53)2 x 7/36 + (4 - 2.53)2 x 5/36 + (5 - 2.53)2 x 3/36 + (6 - 2.53)2 x 1 / 36 = 1.971451

Standard deviation, = sigma = νσ2-V/1.971451 = $ 1.40 mn

Que 3

Expected gain / loss = Expected payoff - cost of the lottery ticket = $ 2,527,778 - 100 = $ 2,527,678

Que 4

Future compounded expected payoff = Future Value of expected payoff Xe = Xe x (1 + R)N where R = interest rate = 5% per annum and N = period = 5 years

Hence, Future compounded expected payoff = 2.53 x (1 + 5%)5 = $ 3.23 mn

Que 5

Yes, there is a correlation between the payoff amount X and the probability associated with it. The probability of the payoff declines as payoff amount increase.

Que 6

As X increases, probability decreases. So, if probability is dependent variable and payoff, X is independent variable then, the depended variable decreases in value as the value of independent variable increase. So, the sign of the regression coefficient must be negative.

Que 7

Regression coefficient is the coefficient of independent variable in the equation of regression. If Y = aX + b is the linear regression equation connecting X and Y,  then Y is the dependent variable and X is the independent variable. Regression coefficient will then be "a", the coefficient of X in the equation of regression. Thus regression coefficient measures the change in the value of dependent variable for a unit change in the value of independent variable. Thus it is also an indicator of sensitivity of dependent variable with respect to independent variable.

Que 8

P value test tells us whether the regression coefficients are significant or not.

The P value in turn is dependent upon t statistics and degrees of freedom.

If P value is less than the common alpha level of 0.05, then the coefficient is statistically significant otherwise it's not.